# Enhancing Virtual Try-On with Synthetic Pairs and Error-Aware Noise Scheduling

### v1

@@ 6.1 Comparison with Existing Methods
Table 2: User studies on VITON-HD. Our EARSB is preferred in an average of 59% cases.
@@
Figure 4: Results on VITON-HD at 5, 10, 25, 50, and 100 sampling steps. Our method consistently improves our baseline starting model GP-VTON (black, dotted line), making it competitive with Stable-VTON (especially at under 50 sampling steps). Legend is shared for all.
Figure 5: Visualizations on VITON-HD (top row) and DressCode (bottom row). Our EARSB+H2G-UH and EARSBSD+H2G-UH better recover the intricate textures in the garment.
@@
Amazon MTurk workers were asked to compare the image quality on two aspects: texture consistency, and image fidelity, comparing our model against GP-GAN and Stable-VTON. We randomly selected 100 pairs from VITON-HD to evaluate on, assigning at least 3 workers per image. Tab. 2 reports the study results, where our method is preferred at least 10% more than the GAN-based GP-VTON and the SD-based Stable-VTON (59% overall).
@@
Diffusion-generated images often show degraded quality with fewer sampling steps. In EARSB, the spatially adaptive noise schedule can preserve the correct clothing textures in the initial image with a low noise level and only fix the erroneous parts, potentially resulting in less image quality degradation with few steps. In Fig. 4, While other SD-based methods have a sharp performance drop with decreasing sampling steps, EARSB and EARSB +H2G-UH show consistent performance across different sampling steps, demonstrating a better trade-off between image quality and computational efficiency.
@@
Fig. 5 gives examples of the generated images using different approaches. The top row is from VITON-HD dataset and the bottom row is from DressCode-Upper. The third images in the two rows are GAN-generated results. We see that our EARSB+H2G-UH/FH in the last column improves the low-quality textures from the GAN-generated images, which are the distorted graphics in the center. More visualized examples can be found in Supp. F.
@@ 6.2 Ablations
Table 3(a) and 3(b) include results of different synthetic data augmentation techniques. In Tab. 3(a), we incorporate H2G-UH into the training of Stable-VTON [19] and CAT-DM [38] on the VITON-HD dataset to validate the effectiveness of synthetic pairs on enhancing existing diffusion methods. Here, we use the number of sampling steps as originally published: 2 for CAT-DM, 50 for Stable-VTON, and 25 for our own EARSB. In Tab. 3(a), +H2G-UH indicates the data augmentation. It shows that training with our synthetic H2G-UH improves most metrics of the above SD-based models.
@@
(a) CAT-DM [38] 8.56 5.90 0.911 0.067
[…]
CAT-DM +H2G-UH 8.36 5.67 0.913 0.063
Stable-VTON [19] 8.25 5.15 0.917 0.056
[…]
Stable-VTON +H2G-UH 8.17 5.04 0.919 0.054
(b) EARSB 8.42 5.25 0.918 0.059
[…]
EARSB + plain H2G-UH 9.64 6.52 0.902 0.073
EARSB pre. H2G-UH 8.35 5.18 0.918 0.059
EARSB +H2G-UH 8.26 5.14 0.919 0.058
(c) Inpaint 9.26 6.33 0.909 0.068
@@
Table 3: Results of ablations on VITON-HD.
Tab. 3(b) explores different ways of incorporating the synthetic data into training. EARSB pre. H2G-UH is pretrained using the synthetic pairs and finetuned on real data, EARSB+plain H2G-UH is trained using the mixed distribution of the real and synthetic pairs without the augmentation label identifying them, and EARSB+H2G-UH uses the mixed data with the identifying label. EARSB pre. H2G-UH is slightly worse than EARSB+H2G-UH. Further, we see the effect of conditioning on the augmentation label, as removing it in EARSB+plain H2G-UH greatly degrades the image quality and causes a significant drop in all metrics.
@@
We explore the importance of the error-aware noise schedule in Tab. 3(c). In our error-aware noise schedule, the error map adapts the noise distribution according to the quality of the image patches in the initial image $x_{1}$. This adaptive approach contrasts with a uniform Gaussian noise application across all locations, which would essentially reduce our model to I2SB. As demonstrated in Tab. 3(c), removing the error map during training (EARSB(w.o. $M$)) results in a substantial decline across all metrics. In another setting, instead of learning the refinement with adaptive noise distribution, we train an inpainting model Inpaint to directly regenerate the erroneous regions in $x_{0}$ using the same model architecture. In this experiment, the image patch in $x_{0}$ is defined to be erroneous if its mean confidence in $M$ is greater than 0.5. In Tab. 3(c), the inpainting model Inpaint shows degraded performance, demonstrating the importance of learning the error-to-ground-truth refinement in our model. Furthermore, we explored using a random error map during the sampling process (EARSB(rand($M$))), which also leads to diminished performance. These results underscore the importance of a meaningful error map in precisely locating and enhancing targeted regions. Additionally, the slight performance degradation observed when removing the classifier guidance (EARSB(w.o. CG)) suggests that the error map employed in our classifier guidance also contributes to overall image quality improvement. Collectively, these findings highlight the crucial role of our adaptive noise schedule in achieving superior results.
@@
Weakly-Supervised Classifier
Our weakly supervised classifier highlights low-quality regions in the initial image $x_{1}$ with only a few hours of labeling. To validate its effectiveness, we train two ablations of our WSC: the Unsupervised Classifier (UC) that only uses image labels (i.e., fake or real), and the Fake/Real Composite Classifier (CC). CC uses both image fake/real labels as well as fake region-level labels which are created by compositing real image patches and fake image patches. The compositing is a fully automatic alternative to manual labeling that provides patch-level labels. We annotated 100 images in the test set to validate their effectiveness. Fig. 6 shows the pixel-level precision-recall curve for retrieving annotated artifact pixels within the bounding boxes using the classifiers’ confidence maps. It is clear that weak supervision remains an incredibly cost-effective approach.
7 Conclusions
This paper proposed an approach that addresses two shortcomings of prior work on virtual try-on. First, we address the limited data availability by introducing a human-to-garment model that generates (human, synthetic garment) pairs from a single image of a clothed individual. Second, we propose a refinement model EARSB that surgically targets localized generation errors from the output of a prior model. EARSB improves the low-quality region of an initially generated image based on a spatially-varying noise schedule that targets known artifacts. Experiments on two benchmark datasets demonstrate that our synthetic data augmentation improves the performance of existing methods and that EARSB enhances the quality of the images generated by prior models.
APPENDIX
@@ Appendix A Implementations Details
For generating the initial image $x_{1}$ in our EARSB training, we employ three try-on GAN models: HR-VTON [21] and SD-VTON [30] and GP-VTON [36]. All human images are processed to maintain their aspect ratio, with the longer side resized to 512 pixels and the shorter side padded with white pixels to reach 512. During training, images undergo random shifting and flipping with a 0.2 probability. The weakly-supervised classifier is trained for 100K iterations with a batch size of 8, while the human-to-garment GAN is trained for 90K iterations with a batch size of 16. EARSB+H2G-UH/FH is trained for 300K iterations with a batch size of 32, incorporating 15% synthetic pairs in each batch. The first 200K iterations are trained on $t\in[0,1]$ while the following 100k iterations are finetuned on $t\in[0,0.5)$ and $t\in[0.5,1]$ respectively following [2]. All models utilize the AdamW optimizer with a learning rate of $10^{-4}$.
@@ Appendix D Ablations on the Quality of the Initial Image
In Tab. 4 we include the FID results of using different try-on GAN models to generate the initial image under the unpaired setting. Baseline means the GAN baseline. As previously stated in Sec. 4.1., we can draw three conclusions from the results: a) our EARSB can refine the GAN-generated image over the GAN baseline; b) the quality of the initial image $x_{1}$ is positively correlated with the quality of the sampled $\hat{x}_{0}$; c) our model achieves higher gains over CAT-DM, which also tries to refine the GAN-generated image but without error-aware noise schedule.
Appendix E Limitations
While our human-to-garment model can effectively generate synthetic paired data for try-on training augmentation, it has some imperfections. The overall quality of synthetic garments is regulated by our filtering criteria (Sec. 3.1.2), yet minor texture deformations occasionally occur. For instance, in Fig. 11, the second pair of the first row shows a misaligned shirt placket in the synthetic garment. This limitation stems partly from the fact that our model is trained in the image domain which lacks 3D information. A potential solution is to utilize DensePose representations extracted from the garment as in [9].
@@ Appendix E Limitations
Appendix F Additional Visualizations
Figures 11 and 12 showcase exemplars from our synthesized datasets H2G-UH and H2G-FH, respectively. We also report quantitative results in Table 5 to evaluate our human-to-garment model on VITON-HD and DressCode-Upper. The generated garment images in Figures 11 and 12 closely mimic the product view of the clothing items, accurately capturing both the shape and texture of the original garments worn by the individuals. This approach to creating synthetic training data for the virtual try-on task is both cost-effective and data-efficient, highlighting the benefits of our proposed human-to-garment model.
Figures 13 and 14 give visualized results of the proposed EARSB and EARSB+H2G-UH. In contrast to previous approaches, EARSB specifically targets and enhances low-quality regions in GAN-generated images, which typically correspond to texture-rich areas. This targeted improvement is evident in the last row of Fig. 13, where EARSB more accurately reconstructs text freinds, and in the third row, where it successfully generates four side buttons. Furthermore, the incorporation of our synthetic dataset H2G-UH with EARSB leads to even more refined details in the generated images, demonstrating the synergistic effect of our combined approach.
Appendix G Ethics
Our work on virtual try-on technology can improve user shopping experiences and reduce clothing returns, which could have positive environmental impacts by reducing waste and transportation emissions. However, we acknowledge several potential ethical considerations:
@@ Appendix G Ethics
  • Environmental impact: While our method aims to reduce environmental impact from returns, the computational resources required for training and running these models have their own carbon footprint. We are committed to optimizing our models for efficiency.

### v3

@@ 5.1 Results
Table 2: User studies on VITON-HD. Our EARSB method is preferred in an average of 59% cases.
@@
Figure 4: Results on VITON-HD at 5, 10, 25, 50, and 100 sampling steps. Our method consistently improves our baseline starting model GP-VTON (black, dotted line), making it competitive with StableVITON (especially at under 50 sampling steps). Legend is shared for all.
Figure 5: Visualizations on VITON-HD (top row) and DressCode (bottom row). Our EARSB+H2G-UH and EARSB(SD)+H2G-UH better recover the intricate textures in the garment.
@@
We asked Amazon MTurk workers to evaluate the texture consistency and image fidelity of synthesized images, comparing our model against GP-VTON and StableVITON. We randomly selected 100 pairs from VITON-HD to evaluate on, assigning at least 3 workers per image. Study results in Tab. 2 report our method is preferred at least 10% more than the GAN-based GP-VTON and the SD-based StableVITON (59% overall).
@@
Diffusion-generated images often show degraded quality with fewer sampling steps. In EARSB, the spatially adaptive noise schedule can preserve correct clothing textures in the initial image with a low noise level and only fix the erroneous parts, potentially resulting in less image quality degradation with fewer steps. In Fig. 4, while other SD-based methods have a sharp performance drop with decreasing sampling steps, EARSB and EARSB+H2G-UH show consistent performance across different sampling steps, demonstrating a better trade-off between image quality and computational efficiency.
@@
Fig. 5 gives examples of generated images from different approaches. The top row is from VITON-HD dataset and the bottom row is from DressCode-Upper. The third images in the two rows are GAN-generated results. We see that our EARSB+H2G-UH/FH in the last column improves the low-quality textures from the GAN-generated images, which are the distorted graphics in the center. Additional visualizations are in Supp. G.
@@ 5.2 Ablations
Tab. 3 incorporates H2G-UH into the training of StableVITON [19] and CAT-DM [39] on the VITON-HD dataset to validate the effectiveness of synthetic pairs on enhancing existing diffusion methods. We observe that that training with our synthetic H2G-UH is effective in improving most metrics.
Data Aug. Unpaired Paired
[…]
CAT-DM [39] None 8.56 5.90 0.911 0.067
CAT-DM H2G-UH 8.36 5.67 0.913 0.063
StableVITON [19] None 8.25 5.15 0.917 0.056
StableVITON H2G-UH 8.17 5.04 0.919 0.054
EARSB(SD) H2G-UH 8.04 4.90 0.925 0.053
Table 3: Comparing the effect of our H2G-UH data augmentation approach on VITON-HD. We bold the best overall results and underline the best results for a single model with and without H2G-UH. Each method uses the number of sampling steps from their paper: 2 for CAT-DM, 50 for StableVITON, and 25 for our own EARSB. We find H2G-UH consistently boosts performance.
Unpaired Paired
[…]
None 8.42 5.25 0.918 0.059
[…]
plain H2G-UH 9.64 6.52 0.902 0.073
pre. H2G-UH 8.35 5.18 0.918 0.059
H2G-UH 8.26 5.14 0.919 0.058
@@
Inpainting 9.26 6.33 0.909 0.068
@@
EARSB (full) 8.42 5.25 0.918 0.059
Table 5: Comparing noise scheduling strategies on VITON-HD. We include a simple inpainting baseline, not using the error map during training (w.o. $M$), a random error map (rand($M$)), and removing classifier guidance (w.o. CG). These results demonstrate the importance of using a meaningful error map.
[…]
@@
Tab. 5 explores the importance of the error-aware noise schedule described in Sec. 4.2, where the error map adapts the noise distribution according to the quality of the image patches in the initial image $x_{1}$. This adaptive approach contrasts with a uniform Gaussian noise application across all locations, which would reduce our model to I2SB. We consider three baselines: Inpainting regenerates rather than refines erroneous regions (where the mean confidence $M$ of an image patch containing an error is greater than 0.5), w.o. $M$ removes the error map during training, and rand($M$) indicates a random error map. As shown in Tab. 5, EARSB outperforms all these baselines, underscoring the importance of a meaningful error map in precisely locating and enhancing targeted regions. Additionally, the slight performance degradation observed when removing the classifier guidance (EARSB(w.o. CG)) suggests that the error map employed in our classifier guidance also contributes to overall image quality improvement. Collectively, these findings highlight the crucial role of our adaptive noise schedule in achieving superior results.
@@
Weakly-Supervised error Classifier (WSC)
Our weakly supervised classifier from Sec. 4.1 highlights low-quality regions in the initial image $x_{1}$ with only a few hours of labeling. To validate its effectiveness, we train two ablations of our WSC: the Unsupervised Classifier (UC) that only uses image labels (i.e., fake or real), and the Fake/Real Composite Classifier (CC). CC uses both image fake/real labels as well as fake region-level labels which are created by compositing real image patches and fake image patches. The compositing is a fully automatic alternative to manual labeling that provides patch-level labels. We annotated 100 images in the test set to validate their effectiveness. Fig. 6 shows the pixel-level precision-recall curve for retrieving annotated artifact pixels within the bounding boxes using the classifiers’ confidence maps. It is clear that weak supervision remains an incredibly cost-effective approach.
6 Conclusion
This paper addresses two shortcomings of prior work on virtual try-on. First, we address the limited data availability by introducing a human-to-garment model that generates (human, synthetic garment) pairs from a single image of a clothed individual. Second, we propose a refinement model EARSB that surgically targets localized generation errors from the output of a prior model. EARSB improves the low-quality region of an initially generated image based on a spatially-varying noise schedule that targets known artifacts. Experiments on two benchmark datasets demonstrate that our synthetic data augmentation improves the performance of prior work and that EARSB enhances the overall image quality.
Acknowledgments This material is based upon work supported, in part, by DARPA under agreement number HR00112020054 and the National Science Foundation under Grant No. DBI-2134696. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the supporting agencies.
@@ Appendix A Implementations Details
For generating the initial image $x_{1}$ in our EARSB training, we employ three try-on GAN models: HR-VTON [21] and SD-VTON [31] and GP-VTON [37]. All human images are processed to maintain their aspect ratio, with the longer side resized to 512 pixels and the shorter side padded with white pixels to reach 512. During training, images undergo random shifting and flipping with a 0.2 probability. The weakly-supervised classifier is trained for 100K iterations with a batch size of 8, while the human-to-garment GAN is trained for 90K iterations with a batch size of 16. As shown in Tab. 6, EARSB+H2G-UH/FH is trained for 300K iterations with a batch size of 32, incorporating 15% synthetic pairs in each batch. The first 200K iterations are trained on $t\in[0,1]$ while the following 100k iterations are finetuned on $t\in[0,0.5)$ and $t\in[0.5,1]$ respectively following [2]. All models utilize the AdamW optimizer with a learning rate of $10^{-4}$.
@@ Appendix D Ablations on the Quality of the Initial Image
In Tab. 7 we include the FID results of using different try-on GAN models to generate the initial image under the unpaired setting. Baseline means the GAN baseline. We can draw three conclusions from the results: a) our EARSB can refine the GAN-generated image over the GAN baseline; b) the quality of the initial image $x_{1}$ is positively correlated with the quality of the sampled $\hat{x}_{0}$; c) our model achieves higher gains over CAT-DM, which also tries to refine the GAN-generated image but without error-aware noise schedule.
Figure 12: Visualization of the generated images in WVTON.
[…]
Stable-VTON [19] 131.76 2.10
EARSB(SD) 127.15 1.67
EARSB(SD) +H2G-UH 120.29 1.18
Table 8: Results on out-of-domain test set WVTON under the unpaired setting. All image background is removed for evaluation.
Appendix E Results on In-the-Wild Dataset
We ran our data-augmented EARSB on the Out-of-Domain test set WVTON [22] under the unpaired setting and removed the image background for evaluation. In Tab. 8, we observe a 7 point gain in FID, showing its good generalization ability. Fig. 12 also shows that EARSB(SD)+H2G-UH better recovers the clothing patterns.
Appendix F Limitations
While our human-to-garment model can effectively generate synthetic paired data for try-on training augmentation, it has some imperfections. The overall quality of synthetic garments is regulated by our filtering criteria (Sec. 3.2), yet minor texture deformations occasionally occur. For instance, in Fig. 13, the second pair of the first row shows a misaligned shirt placket in the synthetic garment. This limitation stems partly from the fact that our model is trained in the image domain which lacks 3D information. A potential solution is to utilize DensePose representations extracted from the garment as in [9].
@@ Appendix F Limitations
Appendix G Additional Visualizations
Figures 13 and 14 showcase exemplars from our synthesized datasets H2G-UH and H2G-FH, respectively. We also report quantitative results in Table 9 to evaluate our human-to-garment model on VITON-HD and DressCode-Upper. The generated garment images in Figures 13 and 14 closely mimic the product view of the clothing items, accurately capturing both the shape and texture of the original garments worn by the individuals. This approach to creating synthetic training data for the virtual try-on task is both cost-effective and data-efficient, highlighting the benefits of our proposed human-to-garment model.
Figures 15 and 16 give visualized results of the proposed EARSB and EARSB+H2G-UH. In contrast to previous approaches, EARSB specifically targets and enhances low-quality regions in GAN-generated images, which typically correspond to texture-rich areas. This targeted improvement is evident in the last row of Fig. 15, where EARSB more accurately reconstructs text freinds, and in the third row, where it successfully generates four side buttons. Furthermore, the incorporation of our synthetic dataset H2G-UH with EARSB leads to even more refined details in the generated images, demonstrating the synergistic effect of our combined approach.
Appendix H Ethics
We acknowledge several potential ethical considerations of our work on virtual try-on: